OpenStudy (anonymous):

What is the geometric mean of 14 and 6

jimthompson5910 (jim_thompson5910):

Hint: the geometric mean of the two numbers x and y is sqrt(x*y)

OpenStudy (anonymous):

do you have to graph this or just find the mean? it would probably be easier to graph it

OpenStudy (anonymous):

Just the mean

OpenStudy (anonymous):

I believe all you do is multiply 14 x 6 and find the square root

jimthompson5910 (jim_thompson5910):

x = 14 y = 6 sqrt(x*y) = sqrt(14*6) = ???